OpenStudy (anonymous):

(sin x)(tan x cos x - cot x cos x) = 1 - 2 cos2x... I can solve it pretty far and then i get stumpped

OpenStudy (anonymous):

i got down to... sinx(sinx)- cos^2x/sinx

OpenStudy (anonymous):

@agent0smith

OpenStudy (psymon):

You're on the right track. Just turn that (sinx - [cos^2(x)/sinx] part into one fraction.

OpenStudy (agent0smith):

so you have \[\Large \frac{ \sin^2 x }{ 1 } - \frac{ \cos^2 x }{ \sin x } = \frac{ \sin^2 x *\sin x}{ 1*\sin x } - \frac{ \cos^2 x }{ \sin x }\]

OpenStudy (anonymous):

no...

OpenStudy (agent0smith):

Yeah, you do... that's the left hand side.

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

sin^3x-cos^2x/(sinx)?

OpenStudy (psymon):

sinx[sinx - [cos^2(x)/sinx]] becomes sinx[(sin^2(x) - cos^2(x))/sinx]. The sinx's then cancel and you go from there.

OpenStudy (agent0smith):

Wait, no you didn't distribute the sinx. I continued from where you were, which is missing a sinx.

OpenStudy (anonymous):

what...what sinx's cancel...why

OpenStudy (anonymous):

where is it missing the sin x :/

OpenStudy (psymon):